# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v1

@@ Abstract
We investigate feature universality in large language models (LLMs), a research field that aims to understand how different models similarly represent concepts in the latent spaces of their intermediate layers. Demonstrating feature universality allows discoveries about latent representations to generalize across several models. However, comparing features across LLMs is challenging due to polysemanticity, in which individual neurons often correspond to multiple features rather than distinct ones. This makes it difficult to disentangle and match features across different models. To address this issue, we employ a method known as dictionary learning by using sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces spanned by neurons corresponding to individual features. After matching feature neurons across models via activation correlation, we apply representational space similarity metrics like Singular Value Canonical Correlation Analysis to analyze these SAE features across different LLMs. Our experiments reveal significant similarities in SAE feature spaces across various LLMs, providing new evidence for feature universality.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks (Bubeck et al., 2023; Naveed et al., 2024; Liu et al., 2023). However, understanding how these models represent and process information internally remains a significant challenge (Bereska & Gavves, 2024). A key question in this domain is whether different LLMs learn similar internal representations, or if each model develops its own unique way of encoding linguistic knowledge. This question of feature universality is crucial for several reasons: it impacts the generalizability of interpretability findings across models, it could accelerate the development of more efficient training techniques, and it may provide insights into safer and more controllable AI systems (Chughtai et al., 2023; Gurnee et al., 2024; Sharkey et al., 2024; Bricken et al., 2023).
@@
Comparing features across LLMs is inherently difficult due to the phenomenon of polysemanticity, where individual neurons often correspond to multiple features rather than distinct ones Elhage et al. (2022). This superposition of features makes it challenging to disentangle and match representations across different models. To address these challenges, we propose a novel approach leveraging sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces. SAEs allow us to decompose the complex, superposed representations within LLMs into distinct features that are easier to analyze and compare across models (Makhzani & Frey, 2013; Cunningham et al., 2023; Bricken et al., 2023). We then apply specific representational space similarity metrics to these SAE features, enabling a rigorous quantitative analysis of feature universality across different LLMs (Klabunde et al., 2023). Our methodology involves several key steps: obtaining SAEs trained on the activations of multiple LLMs to extract interpretable features, developing and applying representational similarity metrics tailored for high-dimensional SAE feature spaces, and conducting extensive experiments comparing SAE features across models with varying architectures, sizes, and training regimes.
@@
To verify the effectiveness of our approach, we present a comprehensive set of experiments and results. Our experimental approach involves solving two issues for representational space similarity measurement: the permutation issue, which involves aligning individual neurons, and the rotational issue, which involves comparing two spaces which have different basis that are rotationally-invariant. Additionally, we present an approach to filter out feature matches with low semantically-meaningful similarity. We demonstrate high correlations between individual SAE features extracted from distinct models, providing evidence for feature universality. Furthermore, we show that semantically meaningful subspaces (e.g., features related to specific concepts, such as Emotions) exhibit even higher similarity across models. We also analyze how feature similarity varies across different layers of LLMs, revealing patterns in how representations evolve through the model depth.
@@ 2 Background
Representational Space Similarity: Representational similarity measures typically compare neural network activations by assessing the similarity between activations from a consistent set of inputs (Klabunde et al., 2023). In this paper, we take a different approach by comparing the representations using the decoder weight matrices $W^{\prime}$ of SAEs, where the columns correspond to the feature neurons. We calculate a similarity score $m(W,W^{\prime})$ for pairs of representations $W$ in SAE $A$ and $W^{\prime}$ from SAE B. We obtain two scores via the following two representational similarity metrics:
Singular Value Canonical Correlation Analysis (SVCCA). Canonical Correlation Analysis (CCA) seeks to identify pairs of canonical variables, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$, from two sets of variables, $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, that are maximally correlated with each other (Hotelling, 1936). Singular Value Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017) enhances CCA by first applying Singular Value Decomposition (SVD) to both $\mathbf{X}$ and $\mathbf{Y}$, which can be expressed as:
@@
Representational Similarity Analysis (RSA). Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) works by first computing, for each space, a Representational Dissimilarity Matrix (RDM) $\mathbf{D}\in\mathbb{R}^{n\times n}$, where each element of the matrix represents the dissimilarity (or similarity) between every pair of data points within a space. The RDM essentially summarizes the pairwise distances between all possible data pairs in the feature space. A common distance metric used is the Euclidean distance. After an RDM is computed for each space, a correlation measurement like Spearman’s rank correlation coefficient is applied to the pair of RDMs to obtain a similarity score. RSA has been applied to measure relational similarity of stimuli across brain regions and of activations across neural network layers (Klabunde et al., 2023).
@@ 3 Methodology
    For the permutation issue: Get mean activation correlation of two decoder SAEs for similar LLMs. Pair each feature A with its max activation correlated feature from model B.
@@
    We find a group of related features in each model by searching for features which activate highly on the top 5 samples’ tokens that belong to a concept category. For example, the concept of "emotions" contains keywords such as "rage". Since the correlations are calculated by tokens, we use single-token keywords. If a keyword from a concept appears at least once, that feature is included. The keywords in each category are given in Appendix D.
@@
Mean Activation Correlation. We take the mean activation correlation between neurons, following the approach of Bricken et al. (2023). We pass a dataset of samples through both models. Then, for each feature in model A, we find its activation correlation with every feature in model B, using tokens as the common instance pairing. For each feature in model A, we find its highest correlated feature in model B. We pair these features together when conducting our experiments. We refer to scores obtained by pairing via mean activation correlation as Paired Scores.
Filter by Top Tokens. We notice that there are many "non-concept mappings" with "non-concept features" that have top 5 token activations on end-of-text / padding tokens, spaces, new lines, and punctuation; these non-concept mappings greatly bring our scores down as they are not accurately pairing features representing concepts, as we describe further in Appendix C. By filtering max activation correlation mappings to remove non-concept mapping features, we significantly raise the similarity scores. We also only keep mappings that share at least one keyword in their top 5 token activations. The list of non-concept keywords is given in the Appendix C.
Filter Many-to-1 Mappings. We find that some of these mappings are many-to-1, which means that more than one feature maps to the same feature. We found that removing many-to-1 mappings slightly increased the scores, and we discuss possible reasons why this may occur in Appendix B. However, the scores still show high similarity even with the inclusion of many-to-1 mappings, as seen in Figures 2. We show only scores of 1-to-1 mappings (filtering out the many-to-1 mappings) in the main text, and show scores of many-to-1 mappings in Appendix B. We use features from the SAEs of Pythia-70m and Gemma-1-2B as "many" features which map onto "one" feature in the SAEs of Pythia-160m and Gemma-2-2B.
@@ 4.1 Experimental Setup
LLM Models. We compare LLMs that use the Transformer model architecture (Vaswani et al., 2017). We compare models that use the same tokenizer because the mean activation correlation pairing relies on comparing two activations using the same tokens. We compare the residual streams of Pythia-70m, which has 6 layers and 512 dimensions, vs Pythia-160m, which has 12 layers and 768 dimensions (Biderman et al., 2023). We compare the residual streams of Gemma-1-2B, which has 18 layers and 2048 dimensions, to Gemma-2-2B, which has 26 layers and 2304 dimensions Team et al. (2024a; b).  [^1] Gemma-2 introduces several architectural improvements over Gemma-1 (Team et al., 2024a; b).
SAE Models. For Pythia, we use pretrained SAEs with 32768 feature neurons trained on residual stream layers from Eleuther (EleutherAI, 2023). For Gemma, we use pretrained SAEs with 16384 feature neurons trained on residual stream layers (Lieberum et al., 2024) We access pretrained Gemma-1-2B SAEs for layers 0, 6, 10, 12, and 17, and Gemma-2-2B SAEs for layers 0 to 26, through the SAELens library (Bloom & Chanin, 2024).
@@ 4.2 SAE Space Similarity
As shown in Figure 9 for Pythia and Figure 10 for Gemma in Appendix E, we also find that mean activation correlation does not always correlate with the global similarity metric scores. For instance, a subset of feature pairings with a moderately high mean activation correlation (eg. 0.6) may yield a low SVCCA score (eg. 0.03).
@@
Pythia-70m vs Pythia-160m . In Figure 5, we compare Pythia-70m, which has 6 layers, to Pythia-160m, which has 12 layers. When compared to Figures 11 and 12 in Appendix E, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. In particular, Figure 5 shows that Layers 2 and 3, which are middle layer of Pythia-70m, are more similar by SVCCA and RSA scores to middle layers 4 to 7 of Pythia-160m compared to other layers. Figure 5 also shows that Layer 5, the last residual stream layer of Pythia-70m, is more similar to later layers of Pythia-160m than to other layers. The number of features after filtering non-concept features and many-to-1 features are given in Tables 5 and 6 in Appendix E.
@@
Gemma-1-2B vs Gemma-2-2B. As shown by the paired SVCCA and RSA results for these models in Figure 7, we find that using SAEs, we can detect highly similar features in layers across Gemma-1-2B and Gemma-2-2B. Compared to mean random pairing scores and p-values in Figures 13 and 14 in Appendix E, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. The number of features after filtering non-concept features and many-to-1 features are given in Tables 7 and 8 in Appendix E.
@@ 4.3 Similarity of Semantically Matched SAE Feature Spaces
Pythia-70m vs Pythia-160m. We compare every layer of Pythia-70m to every layer of Pythia-160m for several concept categories. While many layer pairs have similar semantic subspaces. Middle layers appear to have the semantic subspaces with the highest similarity. Table 1 demonstrates one example of this by comparing the SVCCA score for layer 3 to layer 5 of Pythia-160m, which shows high similarity for semantic subspaces across models, as the concept categories all pass Test 1, having p-values below 0.05. We show the scores for other layer pairs and categories in Figures 15 and 17 in Appendix E, which include RSA scores in Table 9.
@@
Gemma-1-2B vs Gemma-2-2B. In Table 2, we compare L12 of Gemma-1-2B vs L14 of Gemma-2-2B. As shown in Figure 7, this layer pair has a very high similarity for most concept spaces; as such, they likely have high similar semantically-meaningful feature subspaces. Notably, not all concept group are not highly similar; for instance, unlike in Pythia, the Country concept group does not pass the Test 1 as it has a p-value above 0.05. We show the scores for other layer pairs and categories in Figures 19 and 21 in Appendix E.
@@ 5 Related Work
Feature Universality. To the best of our knowledge, only Bricken et al. (2023) has done a quantitative study on individual SAE feature similarity for two 1-layer toy models, finding that individual SAE feature correlations are stronger than individual LLM neuron correlations; however, this study did not analyze the global properties of feature spaces. Gurnee et al. (2024) found evidence of universal neurons across language models by applying pairwise correlation metrics, and taxonimized families of neurons based on patterns of downstream, functional effects. O’Neill et al. (2024) discover "feature families", which represent related hierarchical concepts, in SAEs across two different datasets. Recent work has shown that as vision and language models are trained with more parameters and with better methods, their representational spaces converge to more similar representations (Huh et al., 2024).
@@
Representational Space Similarity. Previous work has studied neuron activation spaces by utilizing metrics to compare the geometric similarities of representational spaces (Raghu et al., 2017; Wang et al., 2018; Kornblith et al., 2019; Klabunde et al., 2024; 2023; Kriegeskorte et al., 2008). It was found that even models with different architectures may share similar representation spaces, hinting at feature universality. However, these techniques have yet to be applied to the feature spaces of sparse autoencoders trained on LLMs. Moreover, these techniques compare the similarity of paired input activations. Our work differs as it compares the similarity of paired feature weights. Essentially, previous works do not show that one can match LLMs by features that both LLMs find in common.
@@
Feature Manifolds. In real world data, features may lie on a manifold, where nearby features respond to similar data. (Olah & Batson, 2023; Bricken et al., 2023). Previous work has discovered the existence of features corresponding to months and days of the week arranged in a circular structure in SAE feature spaces across models (Engels et al., 2024). Studying how feature arrangements generalize across models may shed light on how features lie on a manifold.
@@
Mechanistic Interpretability. Recent work has made notable progress in neuron interpretation (Foote et al., 2023; Garde et al., 2023) and interpreting the interactions between Attention Heads and MLPs (Neo et al., 2024). Other work in mechanistic interpretability has traditionally focused on ’circuits’-style analysis (Elhage et al., 2021), as well as the use of steering vectors to direct models at inference using identified representations (Zou et al., 2023; Turner et al., 2023; Bereska & Gavves, 2024). These approaches have been combined with SAEs to steer models in the more interpretable SAE feature space (Nanda & Conmy, 2024). Steering vectors have been found across models; for instance, Arditi et al. (2024) found "Refusal Vectors" that steered a model to refuse harmful instructions across 13 different models, including Gemma and Llama-2.
@@
AI Safety. Advanced AI systems may not be well aligned with the values of their designers (Ngo et al., 2022), due to phenomena such as goal misgeneralization (Shah et al., 2022; Langosco et al., 2022), deceptive alignment (Hubinger et al., 2024; 2021), sycophancy (Sharma et al., 2023), and reward tampering (Denison et al., 2024). SAEs may be able to find AI safety-related features across models; using SAEs to reliably find universal capabilities, especially in combination with control methods such as steering, may help provide solutions to the above pheonomena.

### v2

@@ Abstract
We investigate feature universality in large language models (LLMs), a research field that aims to understand how different models similarly represent concepts in the latent spaces of their intermediate layers. Demonstrating feature universality allows discoveries about latent representations to generalize across several models. However, comparing features across LLMs is challenging due to polysemanticity, in which individual neurons often correspond to multiple features rather than distinct ones. This makes it difficult to disentangle and match features across different models. To address this issue, we employ a method known as dictionary learning by using sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces spanned by neurons corresponding to individual features. After matching feature neurons across models via activation correlation, we apply representational space similarity metrics on SAE feature spaces across different LLMs. Our experiments reveal significant similarities in SAE feature spaces across various LLMs, providing new evidence for feature universality.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks (Bubeck et al., 2023; Naveed et al., 2024; Liu et al., 2023). However, understanding how these models represent and process information internally remains a significant challenge (Bereska & Gavves, 2024). A key question in this domain is whether different LLMs learn similar internal representations, or if each model develops its own unique way of encoding linguistic knowledge. This question of feature universality is crucial for several reasons: it impacts the generalizability of interpretability findings across models, it could accelerate the development of more efficient training techniques, and it may provide insights into safer and more controllable AI systems (Chughtai et al., 2023b; Gurnee et al., 2024; Sharkey et al., 2024; Bricken et al., 2023).
@@
Comparing features across LLMs is inherently difficult due to the phenomenon of polysemanticity, where individual neurons often correspond to multiple features rather than distinct ones (Elhage et al., 2022). This superposition of features makes it challenging to disentangle and match feature representations across different models. To address these challenges, we propose a novel approach leveraging sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces before matching features. SAEs allow us to decompose the complex, superposed representations within LLMs into features that correspond to distinct semantic concepts that are easier to analyze and compare across models (Makhzani & Frey, 2013; Cunningham et al., 2023; Bricken et al., 2023). We then apply representational space similarity metrics to these SAE features, enabling a rigorous quantitative analysis of feature universality across different LLMs (Klabunde et al., 2023). Our methodology involves several key steps: obtaining SAEs trained on the activations of multiple LLMs to extract interpretable features, developing and applying representational similarity metrics tailored for high-dimensional SAE feature spaces, and conducting extensive experiments comparing SAE features across models with varying architectures, sizes, and training regimes.
@@
To verify the effectiveness of our approach, we present a comprehensive set of experiments and results. Our experimental approach involves solving two issues for representational space similarity measurement: the permutation issue, which involves aligning individual neurons, and the rotational issue, which involves comparing two spaces which have different basis that are rotationally-invariant. Additionally, we present an approach to filter out feature matches with low semantically-meaningful similarity. We demonstrate high correlations between individual SAE features extracted from distinct models, providing evidence for feature universality. Furthermore, we show that semantically meaningful subspaces (e.g., features related to specific concepts, such as Emotions) exhibit even higher similarity across models. We also analyze how feature similarity varies across different layers of LLMs, revealing patterns in how representations evolve through the model depth. Lastly, we examine our results to note the differences which are not captured by SAEs. Thus, our research guides future work to improve upon methods that can better capture universal representations of features, and contributes to ongoing research that aims to improve SAEs (Engels et al., 2024b).
@@ 2 Background
Feature Universality. Previous studies have explored the concept of universality using various definitions Bricken et al. (2023); Chughtai et al. (2023a). In this paper, we examine universality through the lens of two definitions: analogous features (Olah et al., 2020), and representational universality (Gurnee et al., 2024). We define correlated features as features which are correlated in terms of activation similarity, and thus activate highly on the same tokens. We measure the representational universality of feature spaces by comparing their correlated feature pairs. Correlated features that reside in spaces with notable representational similarity are deemed as analogous. Further discussion of these definitions is provided in Appendix §C.
@@
Representational Space Similarity. Representational similarity measures typically compare neural network activations by assessing the similarity between activations from a consistent set of inputs (Klabunde et al., 2023). In this paper, we take a different approach by comparing the representations using the decoder weight matrices $W^{\prime}$ of SAEs, where the columns correspond to the feature neurons. We calculate a similarity score $m(W,W^{\prime})$ for pairs of representations $W$ in SAE $A$ and $W^{\prime}$ from SAE $B$. We obtain two scores via the following two representational similarity metrics:
Singular Value Canonical Correlation Analysis (SVCCA): Canonical Correlation Analysis (CCA) seeks to identify pairs of canonical variables, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$, from two sets of variables, $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, that are maximally correlated with each other (Hotelling, 1936). Singular Value Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017) enhances CCA by first applying Singular Value Decomposition (SVD) to both $\mathbf{X}$ and $\mathbf{Y}$, which can be expressed as:
@@
Representational Similarity Analysis (RSA): Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) works by first computing, for each space, a Representational Dissimilarity Matrix (RDM) $\mathbf{D}\in\mathbb{R}^{n\times n}$, where each element of the matrix represents the dissimilarity (or similarity) between every pair of data points within a space. The RDM essentially summarizes the pairwise distances between all possible data pairs in the feature space. A common distance metric used is the Euclidean distance. After an RDM is computed for each space, a correlation measurement like Spearman’s rank correlation coefficient is applied to the pair of RDMs to obtain a similarity score. RSA has been applied to measure relational similarity of stimuli across brain regions and of activations across neural network layers (Klabunde et al., 2023).
@@ 3 Methodology
    For the permutation issue: Get activation correlations for every feature pair from two decoder SAEs for similar LLMs. Pair each feature from model $A$ with its highest activation correlated feature from model $B$.
@@
    We find a group of related features in each model by searching for features which highly activate on the top 5 samples’ tokens that belong to a concept category. For example, the concept of "emotions" contains keywords such as "rage". Since the correlations are calculated by tokens, we use single-token keywords. If a keyword from a concept appears at least once, that feature is included. The keywords in each category are given in Appendix G.
@@
Highest Activation Correlation. We take the activation correlations between every feature, following the approach of Bricken et al. (2023). We pass a dataset of samples through both models. Then, for each feature in model $A$, we find its activation correlation with every feature in model $B$, using tokens as the common instance pairing. Lastly, for each feature in model $A$, we find its highest correlated feature in model $B$ (as this allows for many-to-1 mappings, this is termed 1A-to-ManyB; the other way around is ManyA-to-1B). We pair these features together when conducting our experiments. We refer to scores obtained by pairing via highest activation correlation as Paired Scores.
Filter by Top Tokens. We notice that there are many "non-concept mappings" with "non-concept features" that have top 5 token activations on end-of-text / padding tokens, spaces, new lines, and punctuation; these non-concept mappings greatly bring our scores down as they are not accurately pairing features representing concepts, as we describe further in Appendix F. By filtering max activation correlation mappings to remove non-concept mapping features, we significantly raise the similarity scores. We also only keep mappings that share at least one keyword in their top 5 token activations. The list of non-concept keywords is given in the Appendix F.
Filter Many-to-1 Mappings. We find that some of these mappings are many-to-1, which means that more than one feature maps to the same feature. We found that removing many-to-1 mappings slightly increased the scores, and we discuss possible reasons why this may occur in Appendix E. However, the scores still show high similarity even with the inclusion of many-to-1 mappings, as seen in Figures 2. We show only scores of 1-to-1 mappings (filtering out the many-to-1 mappings) in the main text, and show scores of many-to-1 mappings in Appendix E. We use features from the SAEs of Pythia-70m and Gemma-1-2B as "many" features which map onto "one" feature in the SAEs of Pythia-160m and Gemma-2-2B.
@@ 4.1 Experimental Setup
LLM Models. We compare LLMs that use the Transformer model architecture (Vaswani et al., 2017). We compare models that use the same tokenizer because the highest activation correlation pairing relies on comparing two activations using the same tokens. We compare the residual streams of Pythia-70m, which has 6 layers and 512 dimensions, to Pythia-160m, which has 12 layers and 768 dimensions (Biderman et al., 2023). We compare the residual streams of Gemma-1-2B, which has 18 layers and 2048 dimensions, to Gemma-2-2B, which has 26 layers and 2304 dimensions (Team et al., 2024a; b).  [^1] Gemma-2 introduces several architectural improvements over Gemma-1 (Team et al., 2024a; b). We compare SAEs trained on the same Tinystories model (Eldan & Li, 2023) in Appendix §A, and show results for across Llama 3 and 3.1 models (Dubey et al., 2024; AI, 2024) in Appendix §B.
SAE Models. For Pythia, we use pretrained SAEs with 32768 feature neurons trained on residual stream layers from Eleuther (EleutherAI, 2023). For Gemma, we use pretrained SAEs with 16384 feature neurons trained on residual stream layers (Lieberum et al., 2024) We access pretrained Gemma-1-2B SAEs for layers 0, 6, 10, 12, and 17, and Gemma-2-2B SAEs for layers 0 to 26, through the SAELens library (Bloom & Chanin, 2024). In Appendix §A, we compare SAEs within model to investigate the extent of feature space similarities on SAEs trained on the same model. In Appendix §B, we compare SAEs trained trained on the layer 25 residual streams of Llama 3-8B and Llama 3.1-8B models.
@@ 4.2 SAE Space Similarity
As shown in Figure 9 for Pythia and Figure 10 for Gemma in Appendix H, we also find that mean of the highest activation correlations does not always correlate with the global similarity metric scores. For instance, a subset of feature pairings with a moderately high mean activation correlation (eg. 0.6) may yield a low SVCCA score (eg. 0.03).
@@
Pythia-70m vs Pythia-160m . In Figure 5, we compare Pythia-70m, which has 6 layers, to Pythia-160m, which has 12 layers. When compared to Figures 11 and 12 in Appendix H, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. In particular, Figure 5 shows that Layers 2 and 3, which are middle layer of Pythia-70m, are more similar by SVCCA and RSA scores to middle layers 4 to 7 of Pythia-160m compared to other layers. Figure 5 also shows that Layer 5, the last residual stream layer of Pythia-70m, is more similar to later layers of Pythia-160m than to other layers. The number of features after filtering non-concept features and many-to-1 features are given in Tables 5 and 6 in Appendix H.
@@
Gemma-1-2B vs Gemma-2-2B. As shown by the paired SVCCA and RSA results for these models in Figure 7, we find that using SAEs, we can detect highly similar features in layers across Gemma-1-2B and Gemma-2-2B. Compared to mean random pairing scores and p-values in Figures 13 and 14 in Appendix H, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. The number of features after filtering non-concept features and many-to-1 features are given in Tables 7 and 8 in Appendix H.
@@ 4.3 Similarity of Semantically Matched SAE Feature Spaces
Pythia-70m vs Pythia-160m. We compare every layer of Pythia-70m to every layer of Pythia-160m for several concept categories. While many layer pairs have similar semantic subspaces. Middle layers appear to have the semantic subspaces with the highest similarity. Table 1 demonstrates one example of this by comparing the SVCCA score for layer 3 to layer 5 of Pythia-160m, which shows high similarity for semantic subspaces across models, as the concept categories all pass Test 1, having p-values below 0.05. We show the scores for other layer pairs and categories in Figures 15 and 17 in Appendix H, which include RSA scores in Table 9.
@@
Gemma-1-2B vs Gemma-2-2B. In Table 2, we compare L12 of Gemma-1-2B vs L14 of Gemma-2-2B. As shown in Figure 7, this layer pair has a very high similarity for most concept spaces; as such, they likely have high similar semantically-meaningful feature subspaces. Notably, not all concept group are not highly similar; for instance, unlike in Pythia, the Country concept group does not pass the Test 1 as it has a p-value above 0.05. We show the scores for other layer pairs and categories in Figures 19 and 21 in Appendix H.
@@ 5 Related Work
Feature Universality. To the best of our knowledge, only Bricken et al. (2023) has done a quantitative study on individual SAE feature similarity for two 1-layer toy models, finding that individual SAE feature correlations are stronger than individual LLM neuron correlations; however, this study did not analyze the global properties of feature spaces. Gurnee et al. (2024) found evidence of universal neurons across language models by applying pairwise correlation metrics, and taxonimized families of neurons based on patterns of downstream, functional effects. O’Neill et al. (2024) discover "feature families", which represent related hierarchical concepts, in SAEs across two different datasets.
Of note is that our paper seeks to measure the extent of universality in terms of analogous features and representational universality, as defined in previous work (Olah et al., 2020; Gurnee et al., 2024), but we do not measure true features, which are exact, atomic hypothesized ground truth features that models may converge towards capturing. This is explained further in Appendix §C. Recent work has shown that as vision and language models are trained with more parameters and with better methods, their representational spaces converge towards more similar representations (Huh et al., 2024). However, we do not claim nor attempt to study in this paper if models converge towards finding the same features.
@@
Representational Space Similarity. Previous work has studied neuron activation spaces by utilizing metrics to compare the geometric similarities of representational spaces (Raghu et al., 2017; Wang et al., 2018b; Kornblith et al., 2019; Klabunde et al., 2024; 2023; Kriegeskorte et al., 2008; Sucholutsky et al., 2023). It was found that even models with different architectures may share similar representation spaces, hinting at feature universality. However, these techniques have yet to be applied to the feature spaces of sparse autoencoders trained on LLMs. Moreover, these techniques compare the similarity of paired input activations. Our work differs as it compares the similarity of paired feature weights. Essentially, previous works do not show that one can match LLMs by features that both LLMs find in common.
@@
Feature Manifolds. In real world data, features may lie on a manifold, where nearby features respond to similar data. (Olah & Batson, 2023; Bricken et al., 2023). Previous work has discovered the existence of features corresponding to months and days of the week arranged in a circular structure in SAE feature spaces across models (Engels et al., 2024a). Studying how feature arrangements generalize across models may shed light on how features lie on a manifold.
@@
Mechanistic Interpretability. Recent work has made notable progress in neuron interpretation (Foote et al., 2023; Garde et al., 2023) and interpreting the interactions between Attention Heads and MLPs (Neo et al., 2024). Other work in mechanistic interpretability has traditionally focused on ’circuits’-style analysis (Elhage et al., 2021), as well as the use of steering vectors to control model behaviors at inference using identified representations (Zou et al., 2023; Turner et al., 2023; Bereska & Gavves, 2024). These approaches have been combined with SAEs to steer models in the more interpretable SAE feature space (Nanda & Conmy, 2024). Steering vectors have been found across models; for instance, Arditi et al. (2024) found "Refusal Vectors" that steered a model to refuse harmful instructions across 13 different models, including Gemma and Llama-2.
@@
AI Safety. Advanced AI systems may not be well aligned with the values of their designers (Ngo et al., 2022), due to phenomena such as goal misgeneralization (Shah et al., 2022; Langosco et al., 2022), deceptive alignment (Hubinger et al., 2024; 2021; Greenblatt et al., 2024), sycophancy (Sharma et al., 2023), and reward tampering (Denison et al., 2024). SAEs may be able to find AI safety-related features across models; using SAEs to reliably find universal capabilities, especially in combination with control methods such as steering Chalnev et al. (2024), may help provide solutions to the above pheonomena.